3. Let f be the function defined by

f (x) =( ex/2)/x.

(a) Is there a y–intercept? (Explain)

(b) Determine the horizontal and vertical asymptotes (if any).

(c) Use the sign pattern for f'(x) to determine

(i) the interval(s) over which f rises and where it falls;

(ii) the local extrema.

(d) Use the sign pattern for f

00 (x) to determine.

(i) where the graph of f is concave up and where it is concave down;

(ii) the inflection point(s) (if any). [20]

4. A rectangular sheet of metal of perimeter 36cm and dimensions x and y is to be rolled into a

cylinder. (Volume of cylinder is V = πr2h.) What values of x and y give the largest volume?

(The thickness of the metal can be neglected.)


1
Expert's answer
2021-05-07T09:51:03-0400

"\displaystyle
(1)\\
f(x) = \frac{e^{\frac{x}{2}}}{x}\\

(a)\\

f(0) = \infty, \\

\textsf{No}\,\, y-\textsf{intercept}\\
(b)\\
\textsf{vertical assumptote does not exist since}\\
e^{x/2} = 0, x \to -\infty\\

f(\infty) = 0\,\,\textsf{Horizontal assumptote is}\,\, y = 0.\\


(c)\\
f'(x) = \frac{1}{2}\frac{e^{\frac{x}{2}}}{x} - \frac{e^{\frac{x}{2}}}{x^2}\\

f(x)\,\,\textsf{rises if}\,\, f'(x) > 0\\
\implies \frac{xe^{\frac{x}{2}} - 2e^{\frac{x}{2}}}{2x^2} > 0\\
\implies e^{\frac{x}{2}}(x - 2) > 0 \\
\implies x > 2\\

f(x)\,\,\textsf{falls}\,\, f'(x) < 0\\
\implies \frac{xe^{\frac{x}{2}} - 2e^{\frac{x}{2}}}{2x^2} < 0\\
\implies e^{\frac{x}{2}}(x - 2) < 0 \\
\implies x < 2\\

\textsf{Local extrema}\,\, 
f(2) = \frac{e}{2}\\

(d)\\
f(x)\,\,\, \textsf{is concave up, iff}\,\,\,f''(x) > 0\\

\implies f(x) = \frac{2x^2e^{\frac{x}{2}} + x^3e^{\frac{x}{2}} - 2x^2e^{\frac{x}{2}} - 4x(x - 2)e^{\frac{x}{2}}}{4x^4} > 0\\
\implies x^2 - 4x + 8 > 0\\
\textsf{The roots of the quadratic}\\
\textsf{equation are complex}\\
\textsf{Hence, the function is not concave up}\\


f(x)\,\,\, \textsf{is concave down, iff}\,\,\,f''(x) < 0\\

\implies f(x) = \frac{2x^2e^{\frac{x}{2}} + x^3e^{\frac{x}{2}} - 2x^2e^{\frac{x}{2}} - 4x(x - 2)e^{\frac{x}{2}}}{4x^4} < 0\\
\implies x^2 - 4x + 8 < 0\\
\textsf{The roots of the quadratic}\\
\textsf{equation are complex}\\
\textsf{Hence, the function is not concave down}\\
\textsf{There is inflexion point iff}\,\,\,f''(x) = 0\\
\textsf{But, since the roots of the}\\
\textsf{quadratic equation are complex,}\\

\textsf{There are no inflexion points}\\


(2)\\\textsf{Area of the rectangle} = xy\\

\textsf{Volume of cylinder} = \pi r^2 h\\

\textsf{A larger surface area gives a}\\
 \textsf{larger volume.}\\

2(x + y) = 36, \, x + y = 18\\

A = x(18 - x)\\
A = x(18 - x)\\
A'(x)= 18 - 2x\\

A \,\, \textsf{is maximum or minimum at}\,\, A'(x) = 0\\

\implies 18 - 2x = 0, x = 9\\

y = 18 - x = 18 - 9 = 9\\

x = 9cm \,\, \textsf{and}\,\, y = 9cm\,\, \textsf{gives the largest volume.}"